What is the magnitude of the intensity at a point which is 0.3 m from a 10 µC charge? (Show your solution)


1
Expert's answer
2021-02-15T17:03:04-0500

Given,

Distance (d) = 0.3m

Charge "(Q)=10 \mu C =10\times 10^{-6}C"

"=10^{-5}C"

Hence, the required electric field "(\overrightarrow{E}) = \frac{Q}{4\pi \epsilon_o d^2}"

Now, substituting the values,

"\Rightarrow \overrightarrow{E}=\frac{10^{-5}\times 9\times 10^{9}}{0.3^2}N/C"


"\Rightarrow \overrightarrow{E}=\frac{9\times 10^4}{0.09}N/C"


"\overrightarrow{E}=1\times 10^6 N/C"